OpenStudy (anonymous):

Rewrite the equation of a line in point-slope form to slope intercept form: y-2=2/3(6x+9)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

y-2=2/3(6x+9) y-2=2/3(6x)+2/3(9) ... Distribute y-2=4x+6 ... Multiply y=4x+6+2 ... Add 2 to both sides. y = 4x + 8 ... Combine like terms.

OpenStudy (anonymous):

thanks for always helping me @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome